# Empirical Bayes for Data integration

### v1

@@ 7. Discussion
Bayesian inference is naturally suited to incorporate previous data or knowledge into the current analysis. Fully Bayesian inference is a powerful framework that is particularly well-suited to this task when one is willing to set a moderately informative prior linking the current and previous data or parameters. In most situations full Bayes enjoys excellent properties. In particular, as $n$ grows its solution often matches that of empirical Bayes and enjoys a frequentist interpretation as conditional probabilities on parameters given the data, under repeated sampling of both data and parameters. There are situations when full and empirical Bayes do not match asymptotically, these are somewhat extreme settings where the prior plays a critical role. In our data integration context, this occurs for example when parameters are divided into two blocks and one is much sparser than the other. From a theoretical viewpoint, it is interesting that the mathematical conditions for model selection consistency are milder when integrates data. As future work, it would be interesting to consider more complex models than linear regression, and cases where $\bm{Z}$ contains many meta-covariates. It would also be interesting to compare our results to those obtained when doing data integration with full data. From a practical viewpoint, in most of our examples the benefits of data integration were moderate but non-negligible.
Performing data integration via empirical Bayes has a computational cost, due to adding more layers to the model. We discussed simple strategies that rendered computations feasible in our examples, but more work in this area is needed. Particularly as one considers transfer learning in large models, where data integration seems particularly promising.
@@ Acknowledgments
PRV acknowledges the support of Departament de Recerca i Universitats de la Generalitat de Catalunya and the European Social Fund. DR was funded by grant Consolidación investigadora CNS2022-135963 by the Agencia Española de Investigación (AE), and PID2022-138268NB-I00 by Ministerio de Ciencia e Innovación (MICIN)/Agencia Española de Investigación(AEI)/10.13039/501100011033 /Fondo Europeo de Desarrollo Regional (FEDER).
@@ 12.1. Auxiliary results
In this subsection we give a series of results instrumental to the proofs of Proofs of Theorems 1 and Theorem 3. We start by recalling some notation and introducing new one. Recall, in (1) and (2) we defined for any $\omega_{1},\ldots,\omega_{B}$,
@@
$S_{b}^{I}(\kappa)=\left\{j:z_{j}=b\right\}\setminus\big{(}S_{b}^{L}(\kappa)\cup S_{b}^{S}(\kappa)\big{)}.$
@@
(9) $\mathcal{T}(\kappa)\;:=\;\Big{\{}\bm{\gamma}\,|\,\bm{\gamma}\,=\,S^{L}(\kappa)\,\cup\,R,\;R\in\mathcal{P}\big{(}S^{I}(\kappa)\big{)}\Big{\}},$
where for a set $A$, $\mathcal{P}(A)$ denotes the power set of $A$.
@@ 12.1.1. Distributional properties and tail bounds
12.1.2. Bound on individual expected posterior model probability
@@
$E\left[\pi(\bm{\gamma}\mid\bm{y},\bm{\omega})\right]\leq\int_{0}^{1}P\left(\bigg{[}1+\frac{\pi(\tilde{\bm{\gamma}}\mid\bm{y},\bm{\omega})}{\pi(\bm{\gamma}\mid\bm{y},\bm{\omega})}\bigg{]}^{-1}\geq u\right){\rm d}u.$
@@ 12.1.3. Convergence to a reduced set of models
Assume A1–A2 hold for $\omega_{1},\ldots,\omega_{B}$, $|S^{I}(\kappa)|=O(1)$ and $|S^{S}_{b}(\kappa)|=O(p_{b}-s_{b})$ (cf (8)) for every $b=1,\ldots,B$. Then,
@@
Finally, consider the third condition in Lemma 8. Note that, since $\tilde{\bm{\gamma}}(\bm{\gamma})\in\mathcal{T}(\kappa)$, we have that $S^{L}(\kappa)\subseteq\tilde{\bm{\gamma}}(\bm{\gamma})$. Moreover we have $\bm{\gamma}\cap S^{I}(\kappa)=\tilde{\bm{\gamma}}(\bm{\gamma})\cap S^{I}(\kappa)$. It follows that $\bm{\gamma}\setminus\tilde{\bm{\gamma}}(\bm{\gamma})\subseteq(S^{I}(\kappa)\cup S^{L}(\kappa))^{C}$, that is the elements of $\bm{\gamma}\setminus\tilde{\bm{\gamma}}(\bm{\gamma})$ are either inactive or belong to $S^{S}(\kappa)$. If $\bm{\gamma}\cap S^{S}(\kappa)=\emptyset$ then $\bm{\gamma}\setminus\tilde{\bm{\gamma}}(\bm{\gamma})\subseteq(\bm{\gamma}^{*})^{C}$ and we immediately get $\mu(\Gamma(\tilde{\bm{\gamma}},\bm{\gamma}),\tilde{\bm{\gamma}}(\bm{\gamma}))=o(A_{\tilde{\bm{\gamma}}(\bm{\gamma})})$ because $\bm{\theta}^{*}_{\bm{\gamma}\setminus\tilde{\bm{\gamma}}(\bm{\gamma})}=\mu(\Gamma(\tilde{\bm{\gamma}},\bm{\gamma}),\tilde{\bm{\gamma}}(\bm{\gamma}))=0$. Assume now that $\bm{\gamma}\cap S^{S}(\kappa)\neq\emptyset$. Using that the rightmost component in (26) is nonnegative and Lemma 3, we have
@@
$A_{\tilde{\bm{\gamma}}(\bm{\gamma})}^{*}\;:=\;\nu\sum_{b=1}^{B}|\bm{\gamma}_{b}\setminus\tilde{\bm{\gamma}}(\bm{\gamma})_{b}|\kappa_{b}+\sum_{b=1}^{B}|\tilde{\bm{\gamma}}(\bm{\gamma})_{b}\setminus\bm{\gamma}_{b}|\left(\tfrac{1-\nu}{6}n\phi^{-1}\rho(\bm{X})\min_{i\in S^{L}_{b}(\kappa)}{\theta^{*}_{i}}^{2}-\nu\kappa_{b}\right)$
Recall that, for every $b$, $\tilde{\bm{\gamma}}(\bm{\gamma})_{b}\setminus\bm{\gamma}_{b}\subseteq S^{L}_{b}(\kappa)$. We have then $\min_{i\in\tilde{\bm{\gamma}}(\bm{\gamma})_{b}\setminus\bm{\gamma}_{b}}{\theta^{*}_{i}}^{2}\geq\min_{i\in S^{L}_{b}(\kappa)}{\theta^{*}_{i}}^{2}$, and by (26), $A_{\tilde{\bm{\gamma}}(\bm{\gamma})}\geq A_{\tilde{\bm{\gamma}}(\bm{\gamma})}^{*}$. It follows that for any $\psi\in(0,1)$ and for all $n$ large enough,
@@
To conclude the second part of the proof we lower-bound $\psi A^{*}_{\tilde{\bm{\gamma}}(\bm{\gamma})}=\sum_{b=1}^{B}|\bm{\gamma}_{b}\setminus\tilde{\bm{\gamma}}(\bm{\gamma})_{b}|\psi\nu\kappa_{b}+\sum_{b=1}^{B}|\tilde{\bm{\gamma}}(\bm{\gamma})_{b}\setminus\bm{\gamma}_{b}|\psi\left(\tfrac{1-\nu}{6}n\phi^{-1}\rho(\bm{X})\min_{i\in S^{L}_{b}(\kappa)}{\theta^{*}_{i}}^{2}-\nu\kappa_{b}\right)$. To do this, we obtain a lower bound for $\psi\nu\kappa_{b}$ and for $\psi\bigg{(}\tfrac{1-\nu}{6}n\phi^{-1}\rho(\bm{X})\min_{i\in S^{L}_{b}(\kappa)}{\theta^{*}_{i}}^{2}-\nu\kappa_{b}\bigg{)}$.
The definition of $S^{L}_{b}(\kappa)$ implies there exists some $g^{\prime}_{b}\to\infty$ such that
(31) $\frac{(1-\nu)n\phi^{-1}\rho(\bm{X})}{6}\min_{i\in S^{L}_{b}(\kappa)}{\theta^{*}_{i}}^{2}\,-\,\kappa_{b}\;=\;\log(s_{b})+g^{\prime}_{b}.$
Let $\delta\in(0,1)$ and denote $\bar{\bm{\gamma}}_{b}=\max\big{\{}\frac{2\log(p_{b}-s_{b})}{f_{b}},\frac{2\log(s_{b})}{g^{\prime}_{b}}\big{\}}$, where $f_{b}$ is given in Assumption A2. Take $\psi=\max_{b=1,\ldots,B}\frac{\xi+\delta+\bar{\bm{\gamma}}_{b}}{1+\bar{\bm{\gamma}}_{b}}$ for some $\xi\in(0,1-\delta)$ then $\psi\in(0,1)$ and we have, for every $b=1,\ldots,B$,
@@
(35) $\psi\bigg{(}\tfrac{1-\nu}{6}n\phi^{-1}\rho(\bm{X})\min_{i\in S^{L}_{b}(\kappa)}{\theta^{*}_{i}}^{2}-\nu\kappa_{b}\bigg{)}\;\geq\;\psi\left(\log(s_{b})+g_{b}^{\prime}\right)\;\geq\;\log(s_{b})+\delta\frac{1}{2}g^{\prime}_{b}.$
where the first inequality follows from (31) and the second inequality from (33). In (30), $\psi A^{*}_{\tilde{\bm{\gamma}}(\bm{\gamma})}=\sum_{b=1}^{B}|\bm{\gamma}_{b}\setminus\tilde{\bm{\gamma}}(\bm{\gamma})_{b}|\psi\nu\kappa_{b}+\sum_{b=1}^{B}|\tilde{\bm{\gamma}}(\bm{\gamma})_{b}\setminus\bm{\gamma}_{b}|\psi\left(\tfrac{1-\nu}{6}n\phi^{-1}\rho(\bm{X})\min_{i\in S^{L}_{b}(\kappa)}{\theta^{*}_{i}}^{2}-\nu\kappa_{b}\right)$. Then by (34) and (35), we get
@@
$d_{b}\;=\;e^{1+\log\big{(}1+\frac{|S^{S}_{b}(\kappa)|}{p_{b}-s_{b}}\big{)}-\delta\tfrac{f_{b}}{2}},\qquad h_{b}\;=\;e^{1-\log\big{(}\frac{s_{b}}{|S_{b}^{L}(\kappa)|}\big{)}-\delta\tfrac{g^{\prime}_{b}}{2}}.$
By assumption $|S^{S}_{b}(\kappa)|=O(p_{b}-s_{b})$, and by definition of $|S_{b}^{L}(\kappa)|$ we have $s_{b}\geq|S_{b}^{L}(\kappa)|$. By Assumption A2 and the definition of $|S_{j}^{L}(\kappa)$, we also have $f_{b}\to\infty$ and $g^{\prime}_{b}\to\infty$, and then $\lim_{n\to\infty}=d_{b}=\lim_{n\to\infty}h_{b}=0$. Using the properties of geometric series, for every $j$ we have
@@ 12.1.4. Convergence of the expected sum of posterior inclusion probabilities
Assume A1–A2 hold for $\omega_{1},\ldots,\omega_{B}$, $|S^{I}(\kappa)|=O(1)$ and $|S^{S}_{b}(\kappa)|=O(p_{b}-s_{b})$ for every $j=1,\ldots,b$. Then
@@
$=$ $\frac{|S^{L}_{b}(\kappa)|}{p_{b}}\sum_{\bm{\gamma}\in\mathcal{T}(\kappa)}\pi(\bm{\gamma}\mid\bm{y},\bm{\omega})+\sum_{\bm{\gamma}\in\mathcal{T}(\kappa)}\pi(\bm{\gamma}\mid\bm{y},\bm{\omega})\sum_{j|z_{j}=b,j\not\in S^{L}_{b}(\kappa)}\frac{I(j\in\bm{\gamma}_{b})}{p_{b}}.$
where the last equality follows from $I(j\in\bm{\gamma}_{b})=1$ for all $j\in S^{L}(\kappa)$ when $\bm{\gamma}\in\mathcal{T}(\kappa)$. The righmost term above and $C_{b}$ are nonnegative, then by the linearity of the expectation

### v2

@@ 7. Discussion
Bayesian inference is naturally suited to incorporate previous data or knowledge. Fully Bayesian inference is particularly appealing when one is willing to set a moderately informative prior linking the current and previous data or parameters. In most situations full Bayes enjoys excellent properties. In particular, as $n$ grows its solution often matches that of empirical Bayes and enjoys a frequentist interpretation as conditional probabilities on parameters given the data, under repeated sampling of both data and parameters. There are situations when full and empirical Bayes do not match asymptotically, these are somewhat extreme settings where the prior plays a critical role. In our data integration context, this occurs for example when parameters are divided into two blocks and one is much sparser than the other. In most of our examples the benefits of using empirical Bayes for data integration were moderate but non-negligible. From a theoretical viewpoint, it is interesting that the mathematical conditions for model selection consistency are milder when one integrates data. Such benefits can be obtained with empirical Bayes, but not will full Bayes. Our theory applies to models with Gaussian errors, such as linear and non-linear additive regression, when there is a single meta-covariate that splits covariates into blocks. Our algorithms apply to fully general settings, they only require estimating posterior inclusion probabilities, and that hyper-parameters only feature in prior inclusion probabilities. In fact, our software includes generalized linear and additive models. Interesting future work includes developing theory beyond Gaussian errors and for cases where $\bm{Z}$ contains many meta-covariates, and comparing our results to those obtained when doing data integration with full data. Our current theory still applies to such settings, provided that one discretizes or clusters the covariates into blocks, but because of the potential information loss our current results would not be tight in such settings.
Performing data integration via empirical Bayes has a computational cost, due to adding more layers to the model. We discussed simple strategies that rendered computations feasible in our examples, but more work in this area is needed. In particular, we proposed generic algorithms to regress prior inclusion probabilities on meta-covariates that apply to any probability model, as long as one can evaluate or approximate marginal likelihoods $p(\bm{y}\mid\bm{\gamma})$ for candidate models $\bm{\gamma}$. An interesting extension is to regress the prior dispersion ($g$ in Zellner’s prior) on meta-covariates $\bm{Z}$, to learn how the magnitude of non-zero coefficients varies with $\bm{Z}$. Such an extension would however complicate the algorithms, and they would likely need to be derived on a case-by-case basis, as opposed to the generic algorithms outlined here.
@@ Acknowledgments
PRV acknowledges the support of Departament de Recerca i Universitats de la Generalitat de Catalunya and the European Social Fund. DR was funded by grant Consolidación investigadora CNS2022-135963 by the Agencia Española de Investigación (AE), and PID2022-138268NB-I00 by Ministerio de Ciencia e Innovación (MICIN)/Agencia Española de Investigación(AEI)/10.13039/501100011033 /Fondo Europeo de Desarrollo Regional (FEDER), and an ICREA Academia fellowship from AGAUR (Generalitat de Catalunya).
@@ 11. Proof of Proposition 1
12. Run times and EM algorithm stability
$p=100$ $p=500$ $p=1,000$ $p=2,000$
($s=2.5$) ($s=5.0$) ($s=11.5$) ($s=19.5$)
Beta-Binomial (LL=5,000) 0.4 5.4 23.1 81.8
Empirical Bayes ($L=$ 5,000; M=M=1,000) 0.5 4.9 28.0 115.2
Empirical Bayes ($L=$ 5,000; M=M=100) 0.3 3.2 18.4 74.0
Empirical Bayes ($L=$ 1,000; M=M=100) 0.1 1.1 5.8 22.0
Table 2. Run times (seconds) in an Ubuntu laptop (13th gen Intel core i7, 32Gb RAM) under simulation settings as in Section 6.1 with $n=500$ and $\bm{\omega}=(-4.6,0.5,0)$. $p$ is the total number of covariates, $s$ the average number that truly have an effect, $L$ the number of MCMC iterations, and $M$ the number of $MC$ iterations in the M-step
$p$ $\hat{\bm{\omega}}^{BB,M=1000}-\hat{\bm{\omega}}^{0,M=1000}$ $\hat{\bm{\omega}}^{BB,M=1000}-\hat{\bm{\omega}}^{0,M=100}$
100 0.00006 0.0001 0.0001 0.0001 0.00005 0.0001
500 0.00337 0.0007 0.0004 0.0059 0.00205 0.0013
Table 3. Mean absolute difference in $\hat{\bm{\omega}}$ for Beta-Binomial initialization $\hat{\bm{\omega}}^{BB}$ vs. $\hat{\bm{\omega}}^{(0)}={\bf 0}$, and $M=1000$ vs. $100$ M-step iterations
[…]
Table 3 assesses the effect of $M$ and of the initial hyper-parameter estimate $\hat{\bm{\omega}}^{(0)}$ in the final $\hat{\bm{\omega}}$ returned by the EM algorithm.
@@ 13.1. Auxiliary results
In this subsection we give a series of results instrumental to the proofs of Theorem 1 and Theorem 3. We start by recalling some notation and introducing new one. Recall, in (3) and (4) we defined for any $\omega_{1},\ldots,\omega_{B}$,
@@
$S_{b}^{I}(\kappa)=\left\{j:z_{j}=b,\theta_{j}^{*}\neq 0\right\}\setminus\big(S_{b}^{L}(\kappa)\cup S_{b}^{S}(\kappa)\big),$
@@
(11) $\mathcal{T}(\kappa)\;:=\;\Big\{\bm{\gamma}\,|\,S^{L}(\kappa)\subseteq\gamma\subseteq S^{L}(\kappa)\cup S^{I}(\kappa)\Big\}.$
@@ 13.1.1. Distributional properties and tail bounds
13.1.2. Bound on the expected posterior probability of an individual model
@@
$E\left[\pi(\bm{\gamma}\mid\bm{y},\bm{\omega})\right]\leq\int_{0}^{1}P\left(\bigg[1+\frac{\pi(\tilde{\bm{\gamma}}\mid\bm{y},\bm{\omega})}{\pi(\bm{\gamma}\mid\bm{y},\bm{\omega})}\bigg]^{-1}\geq u\right){\rm d}u,$
@@ 13.1.3. Convergence to a reduced set of models
Assume that A1–A2 hold for $\omega_{1},\ldots,\omega_{B}$, $|S^{I}(\kappa)|=O(1)$ and $|S^{S}_{b}(\kappa)|=O(p_{b}-s_{b})$ (cf (10)) for every $b=1,\ldots,B$. Then,
@@
Finally, consider the third condition in Lemma 8. Note that, since $\tilde{\bm{\gamma}}(\bm{\gamma})\in\mathcal{T}(\kappa)$, we have that $S^{L}(\kappa)\subseteq\tilde{\bm{\gamma}}(\bm{\gamma})$. Moreover, we have $\bm{\gamma}\cap S^{I}(\kappa)=\tilde{\bm{\gamma}}(\bm{\gamma})\cap S^{I}(\kappa)$. It follows that $\bm{\gamma}\setminus\tilde{\bm{\gamma}}(\bm{\gamma})\subseteq(S^{I}(\kappa)\cup S^{L}(\kappa))^{C}$, that is the elements of $\bm{\gamma}\setminus\tilde{\bm{\gamma}}(\bm{\gamma})$ are either inactive or belong to $S^{S}(\kappa)$. If $\bm{\gamma}\cap S^{S}(\kappa)=\emptyset$ then $\bm{\gamma}\setminus\tilde{\bm{\gamma}}(\bm{\gamma})\subseteq(\bm{\gamma}^{*})^{C}$ and we immediately get $\mu(\Gamma(\tilde{\bm{\gamma}},\bm{\gamma}),\tilde{\bm{\gamma}}(\bm{\gamma}))=o(A_{\tilde{\bm{\gamma}}(\bm{\gamma})})$ because $\bm{\theta}^{*}_{\bm{\gamma}\setminus\tilde{\bm{\gamma}}(\bm{\gamma})}=\mu(\Gamma(\tilde{\bm{\gamma}},\bm{\gamma}),\tilde{\bm{\gamma}}(\bm{\gamma}))=0$. Assume now that $\bm{\gamma}\cap S^{S}(\kappa)\neq\emptyset$. Using that the rightmost component in (28) is nonnegative and Lemma 3, we have
@@
$A_{\tilde{\bm{\gamma}}(\bm{\gamma})}^{*}\;:=\;\nu\sum_{b=1}^{B}|\bm{\gamma}_{b}\setminus\tilde{\bm{\gamma}}(\bm{\gamma})_{b}|\kappa_{b}+\sum_{b=1}^{B}|\tilde{\bm{\gamma}}(\bm{\gamma})_{b}\setminus\bm{\gamma}_{b}|\left(\tfrac{1-\nu}{6}n\phi^{-1}\rho(\bm{X})\min_{j\in S^{L}_{b}(\kappa)}{\theta^{*}_{j}}^{2}-\nu\kappa_{b}\right)$
Recall that, for every $b$, $\tilde{\bm{\gamma}}(\bm{\gamma})_{b}\setminus\bm{\gamma}_{b}\subseteq S^{L}_{b}(\kappa)$. We have then $\min_{i\in\tilde{\bm{\gamma}}(\bm{\gamma})_{b}\setminus\bm{\gamma}_{b}}{\theta^{*}_{i}}^{2}\geq\min_{j\in S^{L}_{b}(\kappa)}{\theta^{*}_{j}}^{2}$, and by (28), $A_{\tilde{\bm{\gamma}}(\bm{\gamma})}\geq A_{\tilde{\bm{\gamma}}(\bm{\gamma})}^{*}$. It follows that for any $\psi\in(0,1)$ and for all $n$ large enough,
@@
To conclude the second part of the proof we lower-bound $\psi A^{*}_{\tilde{\bm{\gamma}}(\bm{\gamma})}=\sum_{b=1}^{B}|\bm{\gamma}_{b}\setminus\tilde{\bm{\gamma}}(\bm{\gamma})_{b}|\psi\nu\kappa_{b}+\sum_{b=1}^{B}|\tilde{\bm{\gamma}}(\bm{\gamma})_{b}\setminus\bm{\gamma}_{b}|\psi\left(\tfrac{1-\nu}{6}n\phi^{-1}\rho(\bm{X})\min_{j\in S^{L}_{b}(\kappa)}{\theta^{*}_{j}}^{2}-\nu\kappa_{b}\right)$. To do this, we obtain a lower bound for $\psi\nu\kappa_{b}$ and for $\psi\bigg(\tfrac{1-\nu}{6}n\phi^{-1}\rho(\bm{X})\min_{j\in S^{L}_{b}(\kappa)}{\theta^{*}_{j}}^{2}-\nu\kappa_{b}\bigg)$.
The definition of $S^{L}_{b}(\kappa)$ implies that there exists some $g^{\prime}_{b}\to\infty$ such that
(33) $\frac{(1-\nu)n\phi^{-1}\rho(\bm{X})}{6}\min_{j\in S^{L}_{b}(\kappa)}{\theta^{*}_{j}}^{2}\,-\,\kappa_{b}\;=\;\log(s_{b})+g^{\prime}_{b}.$
Let $\delta\in(0,1)$ and denote $\bar{\bm{\gamma}}_{b}=\max\big\{\frac{2\log(p_{b}-s_{b})}{f_{b}},\frac{2\log(s_{b})}{g^{\prime}_{b}}\big\}$, where $f_{b}$ is given in Assumption A2. Take $\psi=\max_{b=1,\ldots,B}\frac{\xi+\delta+\bar{\bm{\gamma}}_{b}}{1+\bar{\bm{\gamma}}_{b}}$ for some $\xi\in(0,1-\delta)$ then $\psi\in(0,1)$, and we have, for every $b=1,\ldots,B$,
@@
(37) $\psi\bigg(\tfrac{1-\nu}{6}n\phi^{-1}\rho(\bm{X})\min_{j\in S^{L}_{b}(\kappa)}{\theta^{*}_{j}}^{2}-\nu\kappa_{b}\bigg)\;\geq\;\psi\left(\log(s_{b})+g_{b}^{\prime}\right)\;\geq\;\log(s_{b})+\delta\frac{1}{2}g^{\prime}_{b},$
where the first inequality follows from (33) and the second inequality from (35). In (32), $\psi A^{*}_{\tilde{\bm{\gamma}}(\bm{\gamma})}=\sum_{b=1}^{B}|\bm{\gamma}_{b}\setminus\tilde{\bm{\gamma}}(\bm{\gamma})_{b}|\psi\nu\kappa_{b}+\sum_{b=1}^{B}|\tilde{\bm{\gamma}}(\bm{\gamma})_{b}\setminus\bm{\gamma}_{b}|\psi\left(\tfrac{1-\nu}{6}n\phi^{-1}\rho(\bm{X})\min_{j\in S^{L}_{b}(\kappa)}{\theta^{*}_{j}}^{2}-\nu\kappa_{b}\right)$. Then by (36) and (37), we get
@@
(45) $d_{b}\;=\;e^{1+\log\big(1+\frac{|S^{S}_{b}(\kappa)|}{p_{b}-s_{b}}\big)-\delta\tfrac{f_{b}}{2}},\qquad h_{b}\;=\;e^{1-\log\big(\frac{s_{b}}{|S_{b}^{L}(\kappa)|}\big)-\delta\tfrac{g^{\prime}_{b}}{2}}.$
By assumption $|S^{S}_{b}(\kappa)|=O(p_{b}-s_{b})$, and by definition of $|S_{b}^{L}(\kappa)|$ we have $s_{b}\geq|S_{b}^{L}(\kappa)|$. By Assumption A2 and the definition of $|S_{b}^{L}(\kappa)$, we also have $f_{b}\to\infty$ and $g^{\prime}_{b}\to\infty$, and then $\lim_{n\to\infty}=d_{b}=\lim_{n\to\infty}h_{b}=0$. Using the properties of geometric series, for every $j$ we have
@@ 13.1.4. Convergence of the expected sum of posterior inclusion probabilities
Assume that A1–A2 hold for $\omega_{1},\ldots,\omega_{B}$, $|S^{I}(\kappa)|=O(1)$ and $|S^{S}_{b}(\kappa)|=O(p_{b}-s_{b})$ for every $j=1,\ldots,b$. Then
@@
$=$ $\frac{|S^{L}_{b}(\kappa)|}{p_{b}}\sum_{\bm{\gamma}\in\mathcal{T}(\kappa)}\pi(\bm{\gamma}\mid\bm{y},\bm{\omega})+\sum_{\bm{\gamma}\in\mathcal{T}(\kappa)}\pi(\bm{\gamma}\mid\bm{y},\bm{\omega})\sum_{j|z_{j}=b,j\not\in S^{L}_{b}(\kappa)}\frac{I(j\in\bm{\gamma}_{b})}{p_{b}},$
where the last equality follows from $I(j\in\bm{\gamma}_{b})=1$ for all $j\in S^{L}(\kappa)$ when $\bm{\gamma}\in\mathcal{T}(\kappa)$. The rightmost term above and $C_{b}$ are nonnegative, then by the linearity of the expectation